what is i to the 16th power?
What is i to the 255th power?
What is i to the 5692nd power?

Answer by fractalier(1829)

i^16 = i^0 = 1
i^255 = i^3 = -i
i^5692 = i^0 = 1
You divide the exponent by four, and your remainder becomes the actual exponent...we say it is mod 4...the powers of i repeat every four powers...